Three resistors are wired in a series with a 22 volt battery. The resistances are 22.5 ohm, 33.6 ohm, and 9.9ohm. What is the voltage across the 9.9 ohm resistor?
1
Expert's answer
2012-05-22T12:17:18-0400
Three resistors are wired in a series with a 22 volt battery. The resistances are 22.5 ohm, 33.6 ohm, and 9.9ohm. What is the voltage across the 9.9 ohm resistor?

Total resistance: R = 22.5+33.6+9.9 = 66 ohm.

Current: I = 22 volt / 66 ohm = 1/3 A.

Voltage across the 9.9 ohm resistor: U = 1/3 A · 9.9 ohm = 3.3 V.